Fish Tank With Chloroform

You're conducting a science experiment involving a fish tank placed in a small incubator maintained at 25 C. A student accidentally left an open can of chloroform inside the incubator over-night. Determine the concentration of the chloroform in the fish tank the next day. Use Raoult's Law

Given:

Henry's Law Constant for CHCl3, Hg= 3.2 atm M-1

Saturated Vapor Pressure of CHCL3= 26 kPa

Volume of water in fish tank = 0.16 m3

Air Space in incubator excluded fish tank = 0.5 m3

Mass of Chloroform in the can = 20g

Hint: the first question to ask will be whether pure liquid chloroform remain in the system at equilibrium. To find the answer, you can make an assumption and validate or refute the assumption through calculations. Recall that if a gas is at equilibrium with a pure organic phase, the partial pressure of the chemical is its saturated vapor pressure.
Answer #1

>>Answer

>>Given

Apply Henrys Law

Mi = Pi/H

Pi = partial pressure of CHCl4

V = 0.16 m3 = 0.16*10^3 L = 160 Liters

Vair = 0.5 m3 = 500 L

mass of Chloroform = 20 g

mol of Chloroform = mass/MW = 20/119.38 = 0.16753 mol of CHCl3

Partial Pressure =26 kPa = 26000 Pa --> 0.2566 atm

then...

Mi = Pi/Hi

Mi = (0.2566 atm) / (3.2 atm/M) = 0.0801 M of gas

concentration in tfish tank:

[CH3Cl] in tank = 0.0801 mol per liter

.......................Thank You..............................